OpenStudy (anonymous):

If h(x) = 6 – x, what is the value of (h*h)(10)

ganeshie8 (ganeshie8):

Hint : (h*h)(10) = h(10) * h(10)

OpenStudy (anonymous):

6-x(10) * 6-x(10)??

OpenStudy (usukidoll):

ah this isn't composite functions.. ok I got confused for a second since we have h(x) = 6-x and (h(x))(h(x)) we just multiply (6-x)(6-x) and then evaluate the result when x = 10

OpenStudy (anonymous):

yay 16

OpenStudy (usukidoll):

or we could obtain the result a bit faster if we let x = 10 immediately and have (6-10)(6-10) oh cool you got it! :D

OpenStudy (usukidoll):

(-4)(-4) = 16

OpenStudy (anonymous):

thank you!

OpenStudy (usukidoll):

\[36-6x-6x+x^2 \rightarrow 36-12x+x^2 \] at x = 10 \[36-12(10)+(10)^2 = 36-120+100=16 \] yay!